# Mirror partner for a Klein quartic polynomial

### v2

@@ 1. Introduction
Mirror symmetry, in one of its formulations, relates a complex variety $X$ to a so–called Landau–Ginzburg orbifold $(f,G)$ ([IV90, V89, W93]). In the latter $f$ defines a quasihomogeneous isolated singularity and $G$ is its symmetry group. In particular ${f\in{\mathbb{C}}[x_{1},\dots,x_{N}]}$ is a polynomial, whose partial derivatives all vanish simultaneously if and only if $x_{1}=\dots=x_{N}=0$ and there is a set of natural numbers $d_{f},d_{1},\dots,d_{N}$, s.t. $f(\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N})=\lambda^{d_{f}}f(x_{1},\dots,x_{N})$ for any $\lambda\in{\mathbb{C}}^{\ast}$.
@@
Both $X$ and $(f,G)$ have some essential ${\mathbb{C}}$–algebras associated to them. These are the cohomology ring $H^{*}(X)$ if $X$ is smooth or orbifold cohomology ring $H_{orb}^{*}(X)$ otherwise (cf. [AGV08]). Note that $H^{*}_{orb}(X)=H^{*}(X)$ if $X$ is smooth. To a pair $(f,G)$ one associates the Hochschild cohomology ring ${\mathsf{HH}}^{*}(f,G)$ of the category of $G$–equivariant matrix factorizations ${\mathrm{MF}}_{G}(f)$. The variety $X$ is said to be mirror to the Landau–Ginzburg orbifold $(f,G)$ if there is a ${\mathbb{C}}$–algebra homomorphism ${H_{orb}^{*}(X)\cong{\mathsf{HH}}^{*}(f,G)}$.
@@ 1.1. Landau-Ginzburg – Calabi-Yau mirror symmetry
Then the Berglund–Hübsch–Henningson dual Landau-Ginzburg orbifold $(\widetilde{f},\widetilde{G})$ (see [BH93, BH95, Kreu94]) can be used to construct the mirror variety by
$X_{\widetilde{f},\widetilde{G}}:=\left\{(x_{1},\dots,x_{N})\in{\mathbb{P}}(d_{1},\dots,d_{N})\ |\ \widetilde{f}(x_{1},\dots,x_{N})=0\right\}\Big{/}(\widetilde{G}/\langle j_{\widetilde{f}}\rangle).$
Theorem 14 of [CR11] together with Theorem 2.4 of [K09] give that there is a ${\mathbb{C}}$–vector space isomorphism $\tau:{H_{orb}^{*}(X_{\widetilde{f},\widetilde{G}})\cong{\mathsf{HH}}^{*}(f,G)}$. This suggests calling $X_{\widetilde{f},\widetilde{G}}$ a mirror pair of $(f,G)$. Up to now $\tau$ is proved to be an algebra homomorphism only for the certain examples (cf. [MR11, MS12, BT3]). The major complication here lies in computing the algebra structure of $H_{orb}^{*}(X_{\widetilde{f},\widetilde{G}})$ while the algebra structure of ${\mathsf{HH}}^{*}(f,G)$ was computed explicitly in [BT2, BTW16, BTW17, HLL20, BI22] using the technique developed by D.Shklyarov in [Sh20].
@@ 1.2. Admissible symmetry groups
The algebra $H_{orb}^{*}(X_{\widetilde{f},\widetilde{G}})$ is endowed with a pairing for any $f$. However one can introduce an essential pairing on ${\mathsf{HH}}^{*}(f,G)$ only if $f$ satisfies the Calabi–Yau condition (cf. [PV12]). Generally in mirror symmetry one considers the groups $G$, s.t. $G\subseteq{\mathrm{SL}}(N,{\mathbb{C}})$. These groups are distinguished by the fact that they preserve volume form of ${\mathbb{C}}^{N}$ and therefore the pairing of both algebras.
@@ 1.3. Mirror symmetry for curves
Efimov and Seidel [E12, S11] established an equivalence between the (suitably defined) triangulated categories of a genus $g\geq 2$ surface $S_{g}$ and Landau–Ginzburg orbifold $(f_{g},G)$ with $f_{g}=x_{1}^{2g+1}+x_{2}^{2g+1}+x_{3}^{2g+1}+x_{1}x_{2}x_{3}$. Here the group $G$ is generated by the diagonal operator $\text{diag}(\zeta,\zeta,\zeta^{2g-1})$, s.t. $\zeta=\exp(\frac{2\pi\sqrt{-1}}{2g+1})$. Without giving the details of this mirror theorem we quote the following consequence: there is a ${\mathbb{C}}$–algebra isomorphism $H^{*}(S_{g})\cong{\mathsf{HH}}^{*}(f_{g},G)$ (cf. [E12, S11, Sh20]).
@@ 1.4. Klein quartic
Denote ${\mathrm{SL}}_{f}:={\mathrm{GL}}_{f}\cap{\mathrm{SL}}(N,{\mathbb{C}})$. For any $G\subseteq{\mathrm{SL}}_{f}$ denote by $\widehat{G}$ the central extension of $G$ by ${\mathbb{Z}}/2{\mathbb{Z}}$ obtained by adding the $-\mathrm{id}\in{\mathrm{GL}}_{f}$ to $G$.
@@
  • (c)
    There is a Frobenius algebra isomorphism ${\mathsf{HH}}^{*}(f,\widehat{V}_{4})\cong H^{\ast}(S_{3})$.
Proof.
Proof of the theorem occupies Section 4 along with some preparational propositions given in the preceding sections. ∎
This theorem suggests that the mirror partner of $X_{\widetilde{f},J}$ is $(f,\widehat{V}_{4})$.
Very important consequence of this theorem is the following
Corollary 2.
@@ 2. Notations and terminologies
The polynomial $f$ is said to define an isolated singularity if $\mathrm{Jac}(f)$ is a finite-dimensional. Then set $\mu_{f}:=\dim_{\mathbb{C}}\mathrm{Jac}(f)$ and call it the Milnor number of $f$. For $N=0$ set $\mathrm{Jac}(f):={\mathbb{C}}$ with $\mu_{f}=1$.
@@
Its class is nonzero in $\mathrm{Jac}(f)$ giving the ${\mathbb{C}}$–bilinear nondegenerate pairing $\eta_{f}$ called residue pairing (see [GH94, Chapter 5], [AGV85, Section 5.11]). The value $\eta_{f}([u],[v])$ is taken as the projection of the product $[u][v]$ to the ${\mathbb{C}}$–span of $[\mathrm{hess}(f)]$. In particular, we fix
@@
The ${\mathbb{C}}$–algebra $(A,\circ)$ with a non–degenerate bilinear form $\eta_{A}$ is called a Frobenius algebra if $\eta_{A}(u,v\circ w)=\eta_{A}(u\circ v,w)$ for all $u,v,w\in A$. The pair $(\mathrm{Jac}(f),\eta_{f})$ is an example of a Frobenius algebra.
@@ 2.1. Symmetries
Let $\widetilde{x}_{1},\dots,\widetilde{x}_{N_{g}},\widetilde{x}_{N_{g}+1},\dots,\widetilde{x}_{N}$ be the coordinates of ${\mathbb{C}}^{N}$, s.t. $\widetilde{x}_{1},\dots,\widetilde{x}_{N_{g}}$ are dual to the eigenspace $1$ basis vectors of $g$. In this coordinates we have $f^{g}=f\mid_{\widetilde{x}_{N_{g}+1}=\dots=\widetilde{x}_{N}=0}$. The proof follows now by the same argument as in Proposition 5 of [ET13]. ∎
@@
3. Phase space
This section is devoted to the definition of ${\mathsf{HH}}^{*}(f,G)$ and also some computational aspects of it.
3.1. The total space
@@
Each direct summand $\mathrm{Jac}(f^{g})\xi_{g}$ will be called the $g$–th sector and denoted by ${\mathcal{B}}_{g}^{\prime}$. Any $h\in{\mathrm{GL}}_{f}$ induces a map
@@
It’s much more complicated to define the algebra structure of ${\mathsf{HH}}^{*}(f,G)$ and we don’t do it here because it will not be needed later on. We will only need that the product $\circ$ of ${\mathsf{HH}}^{*}(f,G)$ respects the $G$–grading: $\circ:{\mathcal{B}}_{g}\otimes{\mathcal{B}}_{h}\to{\mathcal{B}}_{gh}$, its unit is $1\xi_{\mathrm{id}}\in{\mathcal{B}}_{\mathrm{id}}$ and $\circ:{\mathcal{B}}_{\mathrm{id}}\otimes{\mathcal{B}}_{\mathrm{id}}\to{\mathcal{B}}_{\mathrm{id}}$ coincides with the product of $(\mathrm{Jac}(f))^{G}$.
@@ 3.2. Computational statements
Let $g\in{\mathrm{GL}}_{f}$ then $g^{\ast}\left(\xi_{g^{k}}\right)=\det g\cdot\xi_{g^{k}}$ for any $k$, s.t. $g^{k}\neq\mathrm{id}$.
@@
This is immediate. ∎
@@
Let $G\subset{\mathrm{GL}}_{f}$ be s.t. $\det g=\pm 1$ for any $g\in G$. Consider the polynomials $\phi_{1}({\bf x}),\phi_{2}({\bf x})\in{\mathbb{C}}[{\bf x}]$, s.t. $\eta_{f}([\phi_{1}],[\phi_{2}])\neq 0$.
Then if $[\phi_{1}]\neq 0$ in $(\mathrm{Jac}(f))^{G}$, then $[\phi_{2}]\neq 0$ in $(\mathrm{Jac}(f))^{G}$ as well.
@@
The Jacobian ideal of $f$ is preserved under the $G$–action. Hence under the assumption of the proposition we have
$[\phi_{1}({\bf x})]\cdot[\phi_{2}({\bf x})]=\alpha[\mathrm{hess}(f)]\neq 0\in(\mathrm{Jac}(f))^{G}$
@@
3.3. $\mathrm{Jac}(f,G)$ for a Klein quartic
@@ 3.3.1. Consider $g\in G\backslash\{\mathrm{id}\}$ with $\det g=1$
The fixed locus of $g$ can be either 00 of $1$–dimensional. If $\mathrm{Fix}(g)=0$, any element $h\in Z(g)$ acts on $\xi_{g}$ by multiplication by $\det h$. In particular, $\xi_{g}$ is $G$–invartiant if $G\subset{\mathrm{SL}}(3,{\mathbb{C}})$ and is not $G$–invariant if $-\mathrm{id}\in G$.
If $\mathrm{Fix}(g)\cong{\mathbb{C}}$, we have $f^{g}=\widetilde{x}^{4}$ where $\widetilde{x}$ is a coordinate of $\mathrm{Fix}(g)$. This gives ${\mathcal{B}}^{\prime}_{g}\cong{\mathbb{C}}\langle[1],[\widetilde{x}],[\widetilde{x}^{2}]\rangle\xi_{g}$. Let $h\in Z(g)$ act on $\widetilde{x}$ by $\rho\in{\mathbb{C}}^{\ast}$. Namely $h(\widetilde{x})=\rho\widetilde{x}$. Then we have
$h^{\ast}([\widetilde{x}^{k}]\xi_{g})=\frac{\det h}{\rho^{k+1}}\cdot[\widetilde{x}^{k}]\xi_{g}.$
In order to compute the direct sum contribution of $g$ to ${\mathsf{HH}}^{*}(f,G)$ it remains to check the multiples $\rho$ of all the elements of $Z(g)$.
@@ 3.3.2. Consider $g\in G$ with $\det g=-1$
The fixed locus of $g$ can be either 00, $1$–dimensional or $2$–dimensional. The first two cases are treated exactly in the same way as above and we concentrate on the last one. Let $\widetilde{x}_{1},\widetilde{x}_{2}$ be the coordinates on $\mathrm{Fix}(g)$. Then $f^{g}$ is quasihomogeneous with weights $(1/4,1/4)$. After some rescaling of the variables $f^{g}$ can either be $\widetilde{x}_{1}^{4}+\widetilde{x}_{2}^{4}$, $\widetilde{x}_{1}^{3}\widetilde{x}_{2}+\widetilde{x}_{2}^{4}$ or $\widetilde{x}_{1}^{3}\widetilde{x}_{2}+\widetilde{x}_{2}^{3}\widetilde{x}_{1}$.
We have ${\mathcal{B}}^{\prime}_{g}\cong{\mathbb{C}}^{9}$ generated by the classes $[\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]$ for some pairs $(a_{1},a_{2})$. Since $\det g=-1$, we have $g^{\ast}(\xi_{g})=-\xi_{g}$ and $g^{\ast}([\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]\xi_{g})=-[\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]\xi_{g}$.

### v3

@@ 1. Introduction
Mirror symmetry, in one of its formulations, relates a complex variety $X$ to a so–called Landau–Ginzburg orbifold $(f,G)$ ([IV90, V89, W93]). In the latter $f$ defines a quasihomogeneous isolated singularity and $G$ is its symmetry group. In particular $f\in{\mathbb{C}}[x_{1},\dots,x_{N}]$} is a polynomial, whose partial derivatives all vanish simultaneously if and only if $x_{1}=\dots=x_{N}=0$ and there is a set of natural numbers $d_{f},d_{1},\dots,d_{N}$, such that $f(\lambda^{d_{1}}x_{1},\dots,\lambda^{d_{N}}x_{N})=\lambda^{d_{f}}f(x_{1},\dots,x_{N})$ for any $\lambda\in{\mathbb{C}}^{\ast}$.
@@
Both $X$ and $(f,G)$ have some essential $\mathbb{C}$}–algebras associated to them. These are the cohomology ring $H^{*}(X)$ if $X$ is smooth or orbifold cohomology ring $H_{orb}^{*}(X)$ otherwise (cf. [AGV08]). Note that $H^{*}_{orb}(X)=H^{*}(X)$ if $X$ is smooth. To a pair $(f,G)$ one associates the Hochschild cohomology ring ${\mathsf{HH}}^{*}(f,G)$ of the category of GG–equivariant matrix factorizations ${\mathrm{MF}}_{G}(f)$. The variety $X$ is said to be mirror to the Landau–Ginzburg orbifold $(f,G)$ if there is a $\mathbb{C}$}–algebra homomorphism $H_{orb}^{*}(X)\cong{\mathsf{HH}}^{*}(f,G)$}.
@@ 1.1. Landau-Ginzburg – Calabi-Yau mirror symmetry
[…]
@@ 1.2. Admissible symmetry groups
[…]
@@ 1.3. Mirror symmetry for curves
P.Acosta investigated in [A14] the vector space structure of the algebras above for the polynomials $f$ that do not satisfy the Calabi–Yau condition. His result shows that $H^{\ast}(X_{f,G})$ and ${\mathsf{HH}}^{\ast}(f,G)$ are of different dimension if $f$ does not satisfy the Calabi–Yau condition.
A smooth genus $g$ surface $S_{g}$ is a Calabi–Yau manifolds only if $g=1$. However, this is essential to ask what is a mirror Landau–Ginzburg model for $S_{g}$ with $g\geq 2$.
Efimov and Seidel [E12, S11] established an equivalence between the (suitably defined) triangulated categories of a genus $g\geq 2$ surface $S_{g}$ and Landau–Ginzburg orbifold $(f_{g},G)$ with $f_{g}=x_{1}^{2g+1}+x_{2}^{2g+1}+x_{3}^{2g+1}+x_{1}x_{2}x_{3}$. Here the group $G$ is generated by the diagonal operator $\text{diag}(\zeta,\zeta,\zeta^{2g-1})$, such that $\zeta=\exp(\frac{2\pi\sqrt{-1}}{2g+1})$. Without giving the details of this mirror theorem we quote the following consequence: there is a $\mathbb{C}$}–algebra isomorphism $H^{*}(S_{g})\cong{\mathsf{HH}}^{*}(f_{g},G)$ (cf. [E12, S11, Sh20]).
The difference between the approaches of Acosta and Efimov–Seidel is that in the first case the variety $X=X_{f,G}$ is given by a Landau–Ginzburg model $(f,G)$ whereas in the seconds case the Landau–Ginzburg model is constituted by the variety $X=S_{g}$. Another important difference is that $H^{\ast}(S_{g})$ is a ${\mathbb{Z}}/2{\mathbb{Z}}$–commutative algebra while Acosta works in the context of FJRW theory (c.f. [FJR]) that is always commutative.
In this note we investigate the homological mirror symmetry result of Efimov and Seidel from the point of view of Berglund–Hübsch–Henningson duality.
@@ 1.4. Klein quartic
Denote ${\mathrm{SL}}_{f}:={\mathrm{GL}}_{f}\cap{\mathrm{SL}}(N,{\mathbb{C}})$. For any $G\subseteq{\mathrm{SL}}_{f}$ denote by $\widehat{G}$ the central extension of $G$ by ${\mathbb{Z}}/2{\mathbb{Z}}$ obtained by adding the $\mathrm{id}\in{\mathrm{GL}}_{f}$ to $G$.
@@
Proof of the theorem occupies Section 4 along with some preparational propositions given in the preceding sections.
This theorem suggests that the mirror partner of $X_{\widetilde{f},J}$ is $(f,\widehat{V}_{4})$. A very important consequence of this theorem is the following observation.
@@
Comparing our theorem to the mirror theorem of Efimov and Seidel it’s important to note that their Landau–Ginzburg orbifold was suitably chosen, while our investigation is in the setup of the general approach of the Landau–Ginzburg orbifolds given by the invertible polynomials. The polynomial $f_{g}$ of Efimov and Seidel is not invertible whatever $g$ one takes.
@@ 1.5. Acknowledgements
The author is grateful to an anonymous referee for many useful comments that helped to improve the text and also find a mistake in the early version of this paper.
@@ 2. Notations and terminologies
The polynomial $f$ is said to define an isolated singularity if $\mathrm{Jac}(f)$ is a finite-dimensional $\mathbb{C}$}–vector space. In this case set $\mu_{f}:=\dim_{\mathbb{C}}\mathrm{Jac}(f)$ and call it the Milnor number of $f$. For $N=0$ set $\mathrm{Jac}(f):={\mathbb{C}}$ with $\mu_{f}=1$.
@@
Its class is nonzero in $\mathrm{Jac}(f)$ giving the $\mathbb{C}$}–bilinear nondegenerate pairing $\eta_{f}$ called the residue pairing (see [GH94, Chapter 5], [AGV85, Section 5.11]). The value $\eta_{f}([u],[v])$ is taken as the projection of the product $[u][v]$ to the $\mathbb{C}$}–span of $\mathrm{hess}(f)]$. In particular, we fix
@@ 2.1. Symmetries
Let $\widetilde{x}_{1},\dots,\widetilde{x}_{N_{g}},\widetilde{x}_{N_{g}+1},\dots,\widetilde{x}_{N}$ be the coordinates of ${\mathbb{C}}^{N}$, such that $\widetilde{x}_{1},\dots,\widetilde{x}_{N_{g}}$ are dual to the eigenspace $1$ basis vectors of $g$. In this coordinates we have $f^{g}=f\mid_{\widetilde{x}_{N_{g}+1}=\dots=\widetilde{x}_{N}=0}$. The proof follows now by the same argument as in Proposition 5 of [ET13]. ∎
@@
3. Phase spaces
This section is devoted to the definition of the state spaces and also the computational aspects of them. These are the A model state space $H^{\ast}(S_{3})$ and the B model state space ${\mathsf{HH}}^{*}(f,G)$.
3.1. A–model total space
Recall that $H^{\ast}_{orb}(S_{g})=H^{\ast}(S_{g})$} because $S_{g}$ is a smooth manifold. The dimension of $H^{\ast}(S_{g})$ is $2+2g$. This is a ${\mathbb{Z}}/2{\mathbb{Z}}$–graded algebra. In particular, we have $H^{\ast}(S_{g})=H^{even}\oplus H^{odd}$ with
$H^{even}={\mathbb{C}}\cdot\delta\oplus{\mathbb{C}}\cdot\gamma\quad H^{odd}=\bigoplus_{k=1}^{g}({\mathbb{C}}\cdot\alpha_{k}\oplus{\mathbb{C}}\cdot\beta_{k}).$
Here $\delta\in H^{0}(S_{g})$, $\gamma\in H^{2}(S_{g})$ and $\alpha_{k},\beta_{k}\in H^{1}(S_{g})$. We can also assume the generators above to be such that the only non–zero products are
$\alpha_{k}\circ\beta_{k}=-\beta_{k}\circ\alpha_{k}=\gamma,\quad\delta\circ a=a,\quad a\in H^{\ast}(S_{g}).$
Such an algebra has a unique up to a $\mathbb{C}$$^{\ast}$–multiple Frobenius algebra structure, however this is not a subject of the current note.
3.2. B–model total space
@@
Each direct summand $\mathrm{Jac}(f^{g})\xi_{g}$ will be called the gg–th sector and denoted by ${\mathcal{B}}_{g}^{\prime}$. Any $h\in{\mathrm{GL}}_{f}$ induces a map
@@
It’s much more complicated to define the algebra structure of ${\mathsf{HH}}^{*}(f,G)$ and we don’t do it here because it will not be needed later on. We will only need that the product $\circ$ of ${\mathsf{HH}}^{*}(f,G)$ respects the GG–grading: $\circ:{\mathcal{B}}_{g}\otimes{\mathcal{B}}_{h}\to{\mathcal{B}}_{gh}$, its unit is $1\xi_{\mathrm{id}}\in{\mathcal{B}}_{\mathrm{id}}$ and $\circ:{\mathcal{B}}_{\mathrm{id}}\otimes{\mathcal{B}}_{\mathrm{id}}\to{\mathcal{B}}_{\mathrm{id}}$ coincides with the product of $(\mathrm{Jac}(f))^{G}$.
@@ 3.3. Computational statements
Let $g\in{\mathrm{GL}}_{f}$ and $k\in{\mathbb{Z}}$ be such that $\mathrm{Fix}(g^{k})=\mathrm{Fix}(g)$ then $g^{\ast}\left(\xi_{g^{k}}\right)=\det g\cdot\xi_{g^{k}}$.
@@
Let $\widetilde{x}_{1},\dots,\widetilde{x}_{N_{g}}$ be the coordinates dual to the eigenvalue $1$ subspace of $g$. Then $(d\widetilde{x}_{1}\wedge\dots\wedge d\widetilde{x}_{N_{g}})\wedge\xi_{g^{k}}$ is a volume form of ${\mathbb{C}}^{N}$. We get
$\det g\cdot(d\widetilde{x}_{1}\wedge\dots\wedge d\widetilde{x}_{N_{g}})\wedge\xi_{g^{k}}=g\left((d\widetilde{x}_{1}\wedge\dots\wedge d\widetilde{x}_{N_{g}})\wedge\xi_{g^{k}}\right)$
$\quad=\rho_{g,g^{k}}\cdot(d\widetilde{x}_{1}\wedge\dots\wedge d\widetilde{x}_{N_{g}})\wedge\xi_{g^{k}}.$
This gives the first claim.
In order to show the second claim consider Eq. (7). We have ${\mathcal{C}}^{G}=G$ and $Z(g)=G$ for any $g\in G$, what completes the proof. ∎
@@
Let $G\subset{\mathrm{GL}}_{f}$ be such that $\det g=\pm 1$ for any $g\in G$. Consider the polynomials $\phi_{1}({\bf x}),\phi_{2}({\bf x})\in{\mathbb{C}}[{\bf x}]$ satisfying $\eta_{f}([\phi_{1}],[\phi_{2}])\neq 0$.
If $\phi_{1}]\neq 0$ in $(\mathrm{Jac}(f))^{G}$, then $\phi_{2}]\neq 0$ in $(\mathrm{Jac}(f))^{G}$ as well.
@@
The Jacobian ideal of $f$ is preserved under the GG–action. Hence under the assumption of the proposition we have
$\phi_{1}({\bf x})]\cdot[\phi_{2}({\bf x})]=\alpha[\mathrm{hess}(f)]\neq 0\in(\mathrm{Jac}(f))^{G}$
@@
3.4. ${\mathsf{HH}}^{*}(f,G)$ for a Klein quartic
@@ 3.4.1. Consider $g\in G\backslash\{\mathrm{id}\}$ with $\det g=1$
The fixed locus of $g$ corresponds to the eigenvalue $1$ subspace and could be either 0 or 11–dimensional. If $\mathrm{Fix}(g)=0$ the element $\xi_{g}$ is a non–zero multiple of a volume form on ${\mathbb{C}}^{N}$ (see Section 3.2). Then any element $h\in Z(g)$ acts on $\xi_{g}$ by multiplication by $\det h$. In particular, $\xi_{g}$ is GG–invartiant if $G\subset{\mathrm{SL}}(3,{\mathbb{C}})$ and is not GG–invariant if $\mathrm{id}\in G$.
If $\mathrm{Fix}(g)\cong{\mathbb{C}}$, we have $f^{g}=\widetilde{x}^{4}$ where $\widetilde{x}$ is a coordinate of $\mathrm{Fix}(g)$. This gives ${\mathcal{B}}^{\prime}_{g}\cong{\mathbb{C}}\langle[1],[\widetilde{x}],[\widetilde{x}^{2}]\rangle\xi_{g}$. Let $h\in Z(g)$ act on $\widetilde{x}$ by $\rho\in{\mathbb{C}}^{\ast}$. Namely $h(\widetilde{x})=\rho\widetilde{x}$. Then we have $\rho_{h,g}=\det h/\rho$ by the same argument as in proof of Proposition 5. This gives
$h^{\ast}([\widetilde{x}^{k}]\xi_{g})=\frac{\det h}{\rho^{k+1}}\cdot[\widetilde{x}^{k}]\xi_{g}.$ (8)
In order to compute the direct sum contribution of $g$ to ${\mathsf{HH}}^{*}(f,G)$ it remains to check the multiples $\rho$ of all the centralizers $Z(g)$.
@@ 3.4.2. Consider $g\in G$ with $\det g=-1$
The fixed locus of $g$ can be either 0, 11–dimensional or 22–dimensional. The first two cases are treated exactly in the same way as above and we concentrate on the last one. Let $\widetilde{x}_{1},\widetilde{x}_{2}$ be the coordinates on $\mathrm{Fix}(g)$. Then $f^{g}$ is quasihomogeneous with weights $(1/4,1/4)$. After some rescaling of the variables $f^{g}$ can either be $\widetilde{x}_{1}^{4}+\widetilde{x}_{2}^{4}$, $\widetilde{x}_{1}^{3}\widetilde{x}_{2}+\widetilde{x}_{2}^{4}$ or $\widetilde{x}_{1}^{3}\widetilde{x}_{2}+\widetilde{x}_{2}^{3}\widetilde{x}_{1}$.
We have ${\mathcal{B}}^{\prime}_{g}\cong{\mathbb{C}}^{9}$ generated by the classes $\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]$ for some pairs $(a_{1},a_{2})$. Since $\det g=-1$, we have $g^{\ast}(\xi_{g})=-\xi_{g}$ and $g^{\ast}([\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]\xi_{g})=-[\widetilde{x}_{1}^{a_{1}}\widetilde{x}_{2}^{a_{2}}]\xi_{g}$. This shows that $({\mathcal{B}}^{\prime}_{g})^{G}=0$.